# Enhancing Biomedical Knowledge Discovery for Diseases: An Open-Source Framework Applied on Rett Syndrome and Alzheimer’s Disease

### v2

@@ 3.1 LaMEL model
where $\{E_{A},E_{B},E_{D},E_{E},E_{F},E_{G},E_{H}\}\in\mathbb{R}^{d}$ and $E_{C}\in\mathbb{R}^{2d}$, $d$ is the embedding size of BiomedBERT base (768) and BiomedBERT large (1024), ; defines the concatenation, * holds for the element-wise multiplication, $t_{[ent]}$, $t_{[/ent]}$ are the embeddings of the start and end special tokens of the entities, $[t_{E}]$ and $[t_{Inter}]$ are the averaged pooled representation of the entities and the intermediate tokens between the entities respectively.
3.2 LaMReD model
LaMReD provides two variations that differ in information synthesis (Fig. 3), aiming to explore the potential effect of different aggregations (Theodoropoulos and Moens, 2023). LaMReDA utilizes element-wise addition to aggregate the entities’ representations, whereas LaMReDM employs element-wise multiplication. The input text is encoded using BiomedBERT (base or large). Following LM encoding, we construct the relation representation by sampling and aggregating tokens from the input sequence. This step enables the model to capture essential features and contextual information relevant to semantic relation classification. To mitigate the risk of overfitting and enhance model generalization, we incorporate a dropout layer (Srivastava et al., 2014) with a probability of 0.3. The linear classification layer takes the aggregated representation and outputs the predicted label.
Following the paradigm proposed by Baldini Soares et al. (2019) and Hogan et al. (2021), we experiment with various approaches for learning relation representations tailored to the relation detection task to empirically ascertain the effectiveness of each strategy (Fig. 3). The explored types of relation representation R are the following:
@@ 3.2 LaMReD model
where $\{R_{A},R_{B},R_{C},R_{D},R_{E},R_{F},R_{G},R_{H},R_{I},\\ R_{J},R_{K},R_{L},R_{M},R_{N},R_{O},R_{P}\}\in\mathbb{R}^{d}$, d is the embedding size of BiomedBERT base (768) and BiomedBERT large (1024), $f()$ is the aggregation function, element-wise addition for LaMReDA and element-wise multiplication for LaMReDM, $l()$ is a linear projection layer with dimension equal to the embedding size, $t_{[ent]_{1}}$, $t_{[/ent]_{1}}$, $t_{[ent]_{2}}$, and $t_{[/ent]_{2}}$ are the embeddings of the start and end special tokens of the first and second entity and $t_{[CLS]}$ is the representation of the special token [CLS]. We define the averaged pooled representation of the entities and the intermediate tokens between the entities as $[t_{E1}]$, $[t_{E2}]$, and $[t_{Inter}]$ correspondingly. In equations 23 and 24, we utilize the localized context vector $cv$ [^9] Additional information is provided in Appendix D. that utilizes the attention heads to locate relevant context for the entity pair and was introduced in ATLOP (Zhou et al., 2021), a state-of-the-art model in document-level relation extraction.
3.3 Experimental setup
@@
The models are trained for 50 epochs and the best checkpoints are retained based on the performance on the development set, measured using the F1-score. We utilize the Adam (Kingma and Ba, 2014) optimizer with a learning rate of 10-5. The batch size is set to 16. We conduct experiments in two distinct setups. In the multi-class setup, we evaluate performance using micro and macro F1-score, considering four relation types: positive, negative, complex, and no relation. In the binary setup, the objective is the prediction of the presence of relation. LaMEL is specifically designed for the binary setup. We utilize the official splits of ReDReS and ReDAD (Tab. 1) and repeat the experiments 10 times with different seeds. To ensure robustness of results, we also employ a 5-fold cross-validation approach. To explore the cross-disease capabilities of our approach, we train the models using one dataset (e.g., ReDReS) and evaluate on the other (e.g., ReDAD), and vice versa. We utilize the relation representation $R_{A}$ (Eq. 9) for LaMReDA and LaMReDM and the entity representation $E_{A}$ (Eq. 1) for LaMEL. These experiments are repeated 10 times with different seeds, and 15% of the training data is excluded to define the development set [^10] Hardware: single NVIDIA RTX 3090 GPU 24GB..
@@
where $x_{1}$ and $x_{2}$ are the projected representations of the two entities, $y$ is the gold-truth label (1 if the entities are correlated, -1 if they are not), $cos()$ is the cosine similarity in the embedding space, and $m$ is the margin parameter that is set to 0. In the inference step, the threshold to predict the presence of relation based on the cosine similarity of the two entity representations is set to 0.5.
4 Results
Tables 2 and 3 report the F1-scores for LaMEL LaMReDA, and LaMReDM, models on the ReDReS and ReDAD datasets. Each cell (except for cross-disease experiments) displays two values: the average F1-score from 10 runs on the original test set (Tab. 1) and the average F1-score from a 5-fold cross-validation. The models perform well across all relation (A-P) and entity (A-H) representations, showing their ability to learn meaningful representations for the semantic relation task regardless of initial token selection. However, we observe patterns regarding the relation representations. In the binary setup, relation representation $R_{G}$ (Eq. 15) yields strong results for both datasets, suggesting that including the [CLS] token representation might be beneficial. In the multi-class setup, relation representations $R_{L}$ (Eq. 20), $R_{J}$ (Eq. 18), and $R_{O}$ (Eq. 23) are effective for both datasets, indicating that the surrounding context is crucial for the more complex task, as $R_{L}$ and $R_{J}$ include the averaged pooled representation of intermediate tokens between entities, and $R_{O}$ leverages the context vector (Zhou et al., 2021). The intra-model comparison reveals that over-parameterization tends to be useful. Using BiomedBERT large generally results in better performance than the base alternative. The BiomedBERT base shows superior performance mainly only in experiments using the original splits of ReDReS (Tab. 1). LaMEL is highly competitive with LaMReDA and LaMReDM, indicating that learning entity embedding spaces optimized for relation detection is promising. LaMEL achieves the highest performance in the 5-fold setup of ReDAD and the original setup of ReDReS, with F1-scores of 91.03% and 91.25% respectively.
@@ 4 Results
Table 2: LaMEL Results (%) in binary setup (BiomedBERT $\square$: base, $\blacksquare$:large): Each cell (unless cross-disease experiments) shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
The inter-model comparison across the same relation representations indicates that the aggregation function does not significantly impact relation detection tasks. Neither LaMReDA (element-wise addition) nor LaMReDM (element-wise multiplication) show a clear advantage over the other. This suggests that the transformer layers of BiomedBERT and the projection layer $l()$ preceding the aggregation are effectively trained in both models to encode the essential information for relation detection, regardless of the aggregation function used. The cross-disease experiments underscore the robustness of the models in both binary and multi-class setups. This robustness supports transfer learning (Zhuang et al., 2020) in semantic relation detection, extending to other diseases, highlighting the potential for broader applications and research endeavors in knowledge discovery.
Data Type1 Binary setup Multi-class setup
Micro Evaluation Macro Evaluation
LaMReDA LaMReDM LaMReDA LaMReDM LaMReDA LaMReDM
ReDReS A 90.72/89.95 90.74/90.57 90.42/89.15 90.71/89.53 74.49/73.91 73.96/75.01 74.36/73.31 74.35/74.91 74.52/74.5 73.66/74.48 74.3/73.06 72.81/75.07
B 90.4/88.79 90.28/89.54 90.47/89.33 90.06/89.74 74.27/74.14 73.72/74.79 74.26/74.45 73.57/75.34 74.32/74.15 73.65/75.19 74.38/74.31 73.11/75.74
C 90.85/89.69 90.75/89.75 90.51/88.84 89.14/89.16 74.93/72.98 73.54/74.59 74.31/72.71 73.69/73.56 74.96/73.75 73.44/74.74 74.1/72.83 73.49/73.88
D 90.55/89.29 90.93/89.25 90.61/89.47 90.53/88.96 73.61/73.85 73.5/74.36 73.02/74.96 73.5/75.77 73.71/74.54 73.7/74.62 73.24/75.12 73.9/76.24
E 89.57/89.39 89.43/88.89 89.57/89.39 89.43/88.89 73.73/75.67 73.68/74.68 73.73/75.67 73.21/74.68 73.95/74.9 74.01/75.1 73.95/74.9 74.01/75.1
F 90.48/89.09 90.62/89.56 90.41/89.19 90.43/89.94 72.86/74.18 73.82/76.55 73.51/74.07 73.33/75.26 72.62/72.82 73.94/76.66 73.32/74.08 74.5/75.84
G 90.78/89.32 90.76/90.26 90.91/89.82 89.47/89.6 74.33/73.35 73.63/73.59 74.05/75.05 73.22/74.34 74.57/73.78 73.31/74.02 74.21/75.13 73.87/74.8
H 90.91/88.88 90.45/88.98 90.29/88.93 89.99/89.14 74.43/73.68 73.62/74.56 73.59/74.06 73.36/74.71 74.48/73.96 73.65/74.62 73.9/74.12 73.42/75.04
I 90.86/89.07 90.47/89.38 90.62/89.35 89.55/89.18 74.75/73.19 73.3/74.78 74.29/74.14 74/74.28 74.8/73.48 73.26/74.9 73.88/74.6 73.91/74.16
J 89.43/89.23 89.65/89.3 89.53/88.99 89.89/89.43 73.75/76.05 74.28/74.55 73.47/75.04 74.43/74.95 74.05/75.09 75.06/74.97 73.52/75.91 74.7/75.53
K 90.1/89.63 90.05/89.26 89.7/89.18 89.64/89.54 74.43/74.4 74.07/75.22 74.47/75.89 74.38/74.97 74.44/74.8 74.23/75.42 74.3/75.81 74.02/74.67
L 89.6/89.86 89.85/89.95 89.82/88.61 90.33/88.93 73.35/74.27 74.32/75.42 73.68/76.5 73.9/76.15 73.16/74.55 74.04/75.52 73.66/75.08 73.52/76.02
M 90.81/90.27 90.07/89.85 90.01/89.3 89.75/89.73 74.21/74.59 74.29/74.85 73.96/74.94 74.32/74.87 74.03/74.04 74.36/74.68 73.72/75.1 73.95/75.59
N 90.73/89.37 90.6/89.71 90.72/88.77 90.63/89.49 74.55/73.49 73.83/73.47 73.86/74.81 73.38/74.58 74.66/74.81 73.97/74.94 74.13/74.82 73.53/74.76
O 90.9/89.94 90.5/89.35 90.9/89.94 90.5/89.35 73.99/74.51 73.77/73.79 73.99/74.51 73.77/73.79 73.83/74.71 73.62/74.27 73.83/74.71 73.62/74.27
P 89.72/89.8 90.31/90.08 89.13/89.19 90.3/89.92 73.37/75.03 73.87/75.02 73.57/75.24 74.41/75.44 73.48/74.79 74.66/75.18 73.54/74.84 73.52/75.79
CD2 87.42 88.93 87.76 88.1 73.09 75.04 74.15 75.35 73.64 74.94 74.38 75.44
ReDAD A 88.31/90.15 89.55/91.07 87.98/90.37 89.14/89.92 77.64/77.07 79.47/78.07 78.34/76.14 80.21/78.17 77.34/77.21 79.26/77.83 78.44/76.4 80.13/78.39
B 87.82/90.57 89.11/90.52 87.66/88.83 88.64/87.11 77.74/77.56 78.65/78.24 78.61/76.43 78.91/78.26 77.13/77.74 78.72/78.31 77.76/76.55 78.98/77.56
C 88.3/89.64 89.21/87.01 88.11/88.79 89.17/89.87 77.14/76.7 79.67/77.89 78.19/76.41 79.32/77.59 77.08/77.05 79.35/77.92 77.82/76.62 79.26/77.78
D 87.33/88.99 89.82/89.25 88.18/89.61 88.8/90.05 78.28/76.73 79.54/75.64 76.81/76.58 78.68/78.37 78.26/76.8 78.47/76.12 76.67/76.84 78.73/78.87
E 88.03/88.63 89.37/90.91 88.03/88.63 89.37/90.91 77.83/77.45 77.54/78.3 77.83/77.45 77.54/78.3 77.75/77.34 77.69/78.41 77.75/77.34 77.69/78.41
F 87.71/89.54 88.45/89.45 87.87/90.11 88.54/90.99 77.59/76.5 79.74/79.23 76.94/76.75 79.68/77.94 77.35/76.33 79.47/79.31 76.95/76.95 79.21/77.38
G 88.17/90.06 89.83/88.96 88.22/89.75 89.55/90.15 77.83/77.64 79.39/78.61 78.13/77.04 79.09/77.91 77.5/77.69 79.58/78.76 77.88/77.56 78.74/77.73
H 88.01/89.14 88.76/90.78 87.73/88.99 88.99/90.39 77.12/76.36 79.57/78.32 78.11/77.81 79.4/78.13 77.09/76.08 78.76/78.88 78.14/77.81 79.18/78.15
I 87.56/88.64 88.05/89.67 87.86/90.14 89.45/90.13 77.77/76.29 79.23/77.7 78.4/76.08 78.99/78.42 77.11/76.56 79.66/78.31 78.49/76.24 78.78/77.97
J 87.99/89.91 88.89/91.12 87.79/90.5 89.06/89.05 77.69/77.48 78.4/78.92 77.14/78.35 78.4/77.55 77.71/77.57 77.94/78.87 76.59/77.92 78.18/77.43
K 88.36/91.01 89.33/90.94 88.01/90.09 89.05/90.89 78.3/78.24 78.25/76.19 78.54/78.13 77.73/77.5 78.11/78.17 78.49/76.1 78.29/78.6 77.42/77.46
M 88.42/90.07 89.57/90.8 88.12/90.52 89.4/90.46 77/77.31 78.85/78.42 78.02/78.5 77.12/77.66 76.62/77.2 79.66/77.85 78.02/78.38 77.08/77.66
N 87.98/90.71 88.94/90.82 88.08/90.43 89.47/90.68 78.21/77.97 78.92/78.24 78.03/77.63 78.78/78.06 78.07/77.3 77.91/78.39 77.74/77.73 78.6/78.1
P 88.02/89.41 88.86/89.93 88.33/90.26 89.51/87.85 78.43/77.45 79.38/76.67 78.44/77.19 79.12/77.7 78.37/76.25 79.44/77.27 78.04/77.31 78.91/77.76
CD2 88.4 89.33 89.01 89.16 73.69 74.29 72.67 72.81 74.13 74.82 72.91 73.76

### v3

@@ IV-A LaMEL model
where $\{E_{A},E_{B},E_{D},E_{E},E_{F},E_{G},E_{H}\}\in\mathbb{R}^{d}$ and $E_{C}\in\mathbb{R}^{2d}$, $d$ is the embedding size of BiomedBERT base (768) and BiomedBERT large (1024), ; defines the concatenation, * holds for the element-wise multiplication, $t_{[ent]}$, $t_{[/ent]}$ are the embeddings of the start and end special tokens of the entities, $[t_{E}]$ and $[t_{Inter}]$ are the averaged pooled representation of the entities and the intermediate tokens between the entities respectively. For example, in the sentence: ”[CLS] [ent] Amyloid fibrils [/ent] are found in many fatal neurodegenerative diseases such as Alzheimer’s disease, Parkinson’s disease, [ent] type II diabetes [/ent], and prion disease. [SEP]”, the two identified entities are amyloid fibrils and type II diabetes. The averaged pooled representation of the two entities corresponds to $[t_{E}]$ independently and $[t_{Inter}]$ is the averaged pooled representation of the intermediate tokens (”are found in many fatal neurodegenerative diseases such as Alzheimer’s disease, Parkinson’s disease”).
Fig. 5 illustrates the model architectures of the LaMEL model (on the right side). The model encodes the input sequence with the large or base version of BiomedBERT model. Different tokens (marked as ”A-H”) are selected to construct the entity representations. Next, the entity representations pass through a dropout layer and then a linear layer to extract the projected representations. The ”;” symbol indicates the concatenation of tokens, while ”*” represents element-wise multiplication between specific tokens.
IV-B LaMReD model
LaMReD provides two variations that differ in information synthesis (Fig. 5), aiming to explore the potential effects of different aggregations [77]. LaMReDA (Fig. 5) utilizes element-wise addition to aggregate the entities’ representations, whereas LaMReDM (Fig. 5) employs element-wise multiplication. The input text is encoded using BiomedBERT (base or large). Following LM encoding, we construct the relation representation by sampling and aggregating tokens from the input sequence. This step enables the model to capture essential features and contextual information relevant to semantic relation classification. To mitigate the risk of overfitting and enhance model generalization, we incorporate a dropout layer [78] with a probability of 0.3. The linear classification layer uses the aggregated representation and outputs the predicted label.
Following the paradigm proposed by [79] and [80], we experiment with various approaches for learning relation representations tailored to the relation detection task to empirically ascertain the effectiveness of each strategy (Fig. 5). The explored types of relation representation R are as follows:
@@ IV-B LaMReD model
where $\{R_{A},R_{B},R_{C},R_{D},R_{E},R_{F},R_{G},R_{H},R_{I},R_{J},R_{K},R_{L},R_{M},\\ R_{N},R_{O},R_{P}\}\in\mathbb{R}^{d}$, d is the embedding size of BiomedBERT base (768) and BiomedBERT large (1024), $f()$ is the aggregation function, element-wise addition for LaMReDA and element-wise multiplication for LaMReDM, $l()$ is a linear projection layer with dimension equal to the embedding size, $t_{[ent]_{1}}$, $t_{[/ent]_{1}}$, $t_{[ent]_{2}}$, and $t_{[/ent]_{2}}$ are the embeddings of the start and end special tokens of the first and second entity and $t_{[CLS]}$ is the representation of the special token [CLS]. We define the averaged pooled representation of the entities and the intermediate tokens between the entities as $[t_{E1}]$, $[t_{E2}]$, and $[t_{Inter}]$, correspondingly. For example, in the sentence: ”[CLS] [ent] Amyloid fibrils [/ent] are found in many fatal neurodegenerative diseases such as Alzheimer’s disease, Parkinson’s disease, [ent] type II diabetes [/ent], and prion disease. [SEP]”, the two identified entities are amyloid fibrils ($[t_{E1}]$) and type II diabetes ($[t_{E2}]$). The averaged pooled representation of the intermediate tokens (”are found in many fatal neurodegenerative diseases such as Alzheimer’s disease, Parkinson’s disease”) corresponds to $[t_{Inter}]$. The embeddings of the start and end special tokens of the first (Amyloid fibrils) and second (type II diabetes) entities are $t_{[ent]_{1}}$, $t_{[/ent]_{1}}$, $t_{[ent]_{2}}$, and $t_{[/ent]_{2}}$, respectively. In equations (23) and (24), we utilize the localized context vector $cv$ (additional information is provided in Appendix D) which utilizes the attention heads to locate relevant context for the entity pair and was introduced in ATLOP [81], a state-of-the-art model in document-level relation extraction.
Fig. 5 illustrates the model architectures of the LaMReDA and LaMReDM models (on the left side). The models start by encoding the input sequence with the BiomedBERT model, which can be either the large or base version. Specific tokens in the input sequence are designated to define the relation representation (labeled as ”A-P”). These relation representations are passed through a linear projection layer, followed by a dropout layer. Finally, a classification layer performs the relation prediction. The figure uses the symbol ”#” to denote element-wise addition and element-wise multiplication for LaMReDA and LaMReDM, respectively.
IV-C Experimental setup
@@
The models are trained for 50 epochs and the best checkpoints are retained based on the performance on the development set, measured using the F1-score. We utilize the Adam [82] optimizer with a learning rate of 10-5. The batch size is set to 16 (Table III). We conduct experiments in two distinct setups. In the multi-class setup, we evaluate performance using micro and macro F1-score, considering four relation types: positive, negative, complex, and no relation. In the binary setup, the objective is to predict whether a relation exists and we assess performance using the micro F1-score. LaMEL is specifically designed for the binary setup. We utilize the official splits of ReDReS and ReDAD (Table II) and repeat the experiments 10 times with different seeds (random numbers that are selected to initialize various processes and pseudorandom number generators, such as the random shuffle of the train set before the beginning of a training epoch or the CUDA initialization, ensuring that the results are reproducible). The following numbers are used: 42, 3, 7, 21, 77, 24, 69, 96, 44, 11. The repetition of the experiments is a common practice in the research community to ensure the robustness of the results. We also employ a 5-fold cross-validation approach to ensure the robustness of results. To explore the cross-disease capabilities of our approach, we train the models using one dataset (e.g., ReDReS) and evaluate on the other (e.g., ReDAD), and vice versa. We utilize the relation representation $R_{A}$ (9) for LaMReDA and LaMReDM and the entity representation $E_{A}$ (1) for LaMEL. These experiments are repeated 10 times with different seeds, and 15% of the training data is excluded to define the development set. We use a single NVIDIA RTX 3090 GPU of 24GB to execute the experiments.
@@
where $x_{1}$ and $x_{2}$ are the projected representations of the two entities, $y$ is the gold-truth label (1 if the entities are correlated, -1 if they are not), $\mathrm{cos()}$ is the cosine similarity in the embedding space, and $m$ is the margin parameter that is set to 0. In the inference step, the threshold for predicting the presence of relation based on the cosine similarity of the two entity representations is set to 0.5 (Table III).
TABLE III: Summary of the experimental parameters.
Parameter Value
Epochs 50
Learning Rate 10-5
Batch Size 16
Margin Parameter $m$ 0
Inference Threshold 0.5
TABLE IV: LaMReDA and LaMReDM Results (%) in binary and multi-class setup (BiomedBERT $\square$: base,$\blacksquare$: large): Each cell (unless cross-disease experiments) shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
Data Type1 Binary setup Multi-class setup
Micro Evaluation Macro Evaluation
LaMReDA LaMReDM LaMReDA LaMReDM LaMReDA LaMReDM
ReDReS A 90.72/89.95 90.74/90.57 90.42/89.15 90.71/89.53 74.49/73.91 73.96/75.01 74.36/73.31 74.35/74.91 74.52/74.50 73.66/74.48 74.30/73.06 72.81/75.07
B 90.40/88.79 90.28/89.54 90.47/89.33 90.06/89.74 74.27/74.14 73.72/74.79 74.26/74.45 73.57/75.34 74.32/74.15 73.65/75.19 74.38/74.31 73.11/75.74
C 90.85/89.69 90.75/89.75 90.51/88.84 89.14/89.16 74.93/72.98 73.54/74.59 74.31/72.71 73.69/73.56 74.96/73.75 73.44/74.74 74.10/72.83 73.49/73.88
D 90.55/89.29 90.93/89.25 90.61/89.47 90.53/88.96 73.61/73.85 73.50/74.36 73.02/74.96 73.21/75.77 73.71/74.54 73.70/74.62 73.24/75.12 73.90/76.24
E 89.57/89.39 89.43/88.89 89.57/89.39 89.43/88.89 73.73/75.67 73.68/74.68 73.73/75.67 73.68/74.68 73.95/74.90 74.01/75.10 73.95/74.90 74.01/75.10
F 90.48/89.09 90.62/89.56 90.41/89.19 90.43/89.94 72.86/74.18 73.82/76.55 73.51/74.07 73.33/75.26 72.62/72.82 73.94/76.66 73.32/74.08 74.50/75.84
G 90.78/89.32 90.76/90.26 90.49/89.82 89.47/89.60 74.33/73.35 73.63/73.59 74.05/75.05 73.22/74.34 74.57/73.78 73.31/74.02 74.21/75.13 73.87/74.80
H 90.91/88.88 90.45/88.98 90.29/88.93 89.99/89.14 74.43/73.68 73.62/74.56 73.59/74.06 73.36/74.71 74.48/73.96 73.65/74.62 73.90/74.12 73.42/75.04
I 90.86/89.07 90.47/89.38 90.62/89.35 89.55/89.18 74.75/73.19 73.30/74.78 74.29/74.14 74.00/74.28 74.80/73.48 73.26/74.90 73.88/74.60 73.91/74.16
J 89.43/89.23 89.65/89.30 89.53/88.99 89.89/89.43 73.75/76.05 74.28/74.55 73.47/75.04 74.43/74.95 74.05/75.09 75.06/74.97 73.52/75.91 74.70/75.53
K 90.10/89.63 90.05/89.26 89.70/89.18 89.64/89.54 74.43/74.40 74.07/75.22 74.47/75.89 74.38/74.97 74.44/74.80 74.23/75.42 74.30/75.81 74.02/74.67
L 89.60/89.86 89.85/89.95 89.82/88.61 90.33/88.93 73.35/74.27 74.32/75.42 73.68/76.50 73.90/76.15 73.16/74.55 74.04/75.52 73.66/75.08 73.52/76.02
M 90.81/90.27 90.07/89.85 90.01/89.30 89.75/89.73 74.21/74.59 74.29/74.85 73.96/74.94 74.32/74.87 74.03/74.04 74.36/74.68 73.72/75.10 73.95/75.59
N 90.73/89.37 90.60/89.71 90.72/88.77 90.63/89.49 74.55/73.49 73.83/73.47 73.86/74.81 73.38/74.58 74.66/74.81 73.97/74.94 74.13/74.82 73.53/74.76
O 90.90/89.94 90.50/89.35 90.90/89.94 90.50/89.35 73.99/74.51 73.77/73.79 73.99/74.51 73.77/73.79 73.83/74.71 73.62/74.27 73.83/74.71 73.62/74.27
P 89.72/89.80 90.31/90.08 89.13/89.19 90.30/89.92 73.37/75.03 73.87/75.02 73.57/75.24 74.41/75.44 73.48/74.79 74.66/75.18 73.54/74.84 73.52/75.79
CD2 87.42 88.93 87.76 88.10 73.09 75.04 74.15 75.35 73.64 74.94 74.38 75.44
ReDAD A 88.31/90.15 89.55/91.07 87.98/90.37 89.14/89.92 77.64/77.07 79.47/78.07 78.34/76.14 80.21/78.17 77.34/77.21 79.26/77.83 78.44/76.40 80.13/78.39
B 87.82/90.57 89.11/90.52 87.66/88.83 88.64/87.11 77.74/77.56 78.65/78.24 78.61/76.43 78.91/78.26 77.13/77.74 78.58/78.31 77.76/76.55 78.98/77.56
C 88.30/89.64 89.21/87.01 88.11/88.79 89.17/89.87 77.14/76.70 79.67/77.89 78.19/76.41 79.32/77.59 77.08/77.05 79.35/77.92 77.82/76.62 79.26/77.78
D 87.33/88.99 89.82/89.25 88.18/89.61 88.80/90.05 78.28/76.73 79.54/75.64 76.81/76.58 78.68/78.37 78.26/76.80 78.47/76.12 76.67/76.84 78.73/78.87
E 88.03/88.63 89.37/90.91 88.03/88.63 89.37/90.91 77.83/77.45 77.54/78.30 77.83/77.45 77.54/78.30 77.75/77.34 77.44/78.41 77.75/77.34 77.44/78.41
F 87.71/89.54 88.45/89.45 87.87/90.11 88.54/90.99 77.59/76.50 79.74/79.23 76.94/76.75 79.68/77.94 77.35/76.33 79.19/79.31 76.95/76.95 79.21/77.38
G 88.17/90.06 89.83/88.96 88.22/89.75 89.55/90.15 77.83/77.64 79.39/78.61 78.13/77.04 79.09/77.91 77.50/77.69 79.12/78.76 77.88/77.56 78.74/77.73
H 88.01/89.14 88.76/90.78 87.73/88.99 88.99/90.39 77.12/76.36 79.57/78.32 78.11/77.81 79.40/78.13 77.09/76.08 78.65/78.88 78.14/77.81 79.18/78.15
I 87.56/88.64 88.05/89.67 87.86/90.14 89.45/90.13 77.77/76.29 79.23/77.70 78.40/76.08 78.99/78.42 77.11/76.56 79.05/78.31 78.49/76.24 78.78/77.97
J 87.99/89.91 88.89/91.12 87.79/90.50 89.06/89.05 77.69/77.48 78.40/78.92 77.14/78.35 78.40/77.55 77.71/77.57 78.57/78.87 76.59/77.92 78.18/77.43
K 88.36/91.01 89.33/90.94 88.01/90.09 89.05/90.89 78.30/78.24 78.25/76.19 78.54/78.13 77.73/77.50 78.11/78.17 78.49/76.10 78.29/78.60 77.42/77.46
M 88.42/90.07 89.57/90.80 88.12/90.52 89.40/90.46 77.00/77.31 78.85/78.42 78.02/78.50 77.12/77.66 76.62/77.20 79.66/77.85 78.02/78.38 77.08/77.66
N 87.98/90.71 88.94/90.82 88.08/90.43 89.47/90.68 78.21/77.97 78.92/78.24 78.03/77.63 78.78/78.06 78.07/77.30 77.91/78.39 77.74/77.73 78.60/78.10
P 88.02/89.41 88.86/89.93 88.33/90.26 89.51/87.85 78.43/77.45 79.38/76.67 78.44/77.19 79.12/77.70 78.37/76.25 79.44/77.27 78.04/77.31 78.91/77.76
CD2 88.40 89.33 89.01 89.16 73.69 74.29 72.67 72.81 74.13 74.82 72.91 73.76
  • 1
    Type of Relation Representation.
  • 2
IV-D Experiments with Additional Language Models
Assessing the impact of different language models on the performance of the LaMReDA and LaMEL models, we conduct experiments using additional models to replace the BioMedBERT model (Fig. 5). Specifically, we evaluate the effect of BioLinkBERT [83] (base model) and BioGPT [84], both of which are pre-trained on PubMed abstracts. BioLinkBERT is selected to determine whether leveraging the links between documents in the pre-training corpus benefits the relation detection task of our study. This model uses citations between PubMed articles to model document connections, which may enhance its ability to capture semantic relations. BioGPT, as a GPT-based [85, 86] transformer decoder [10] representative, is incorporated to explore the generative model’s capabilities in capturing semantic relations between entities. [83] state that the pre-training data of BioLinkBERT and BioMedBERT are identical [^5] https://pubmed.ncbi.nlm.nih.gov. Papers published before Feb. 2020., except that they also use the citations between PubMed articles. [84] clarify that BioGPT’s pre-training includes PubMed abstracts updated before 2021. Hence, all language models in our study are pre-trained on identical or nearly identical datasets, making the comparison between them reliable.
We conduct experiments using a variety of representations, specifically testing the $R_{A}$ (9), $R_{D}$ (12), $R_{E}$ (13), $R_{F}$ (14), $R_{G}$ (15), and $R_{O}$ (23) relation representations of LaMReDA, as well as the $E_{A}$ (1), $E_{D}$ (4), $E_{E}$ (5), and $E_{H}$ (8)) entity representations of LaMEL. The experimental setup follows the same procedure outlined in Subsection IV-C (Table III), and the experiments are performed on both the original split (Table II) and in a 5-fold cross-validation setting.
TABLE V: LaMReDA Results (%) in binary and multi-class setup utilizing BiomedBERT-base, BioLinkBERT-base, and BioGPT as backbone language model: Each cell shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
Data Type1 Binary setup Multi-class setup
Micro Evaluation Macro Evaluation
ReDReS A 90.72/89.95 91.15/89.90 90.15/88.83 74.49/73.91 74.06/75.13 72.22/72.65 74.52/74.50 74.11/75.09 71.80/72.95
E 89.57/89.39 89.98/90.28 89.49/87.19 73.73/75.67 73.22/74.40 70.17/70.98 73.95/74.90 73.38/74.97 69.97/70.69
F 90.48/89.09 90.72/89.84 89.68/88.35 72.86/74.18 74.22/75.12 71.36/73.57 72.62/72.82 74.13/75.10 72.00/73.82
G 90.78/89.32 90.57/89.57 89.93/88.41 74.33/73.35 74.00/75.65 72.68/73.14 74.57/73.78 74.32/75.94 72.48/73.43
O 90.90/89.94 90.75/90.12 89.51/87.82 73.99/74.51 74.00/73.65 66.48/68.67 73.83/74.71 73.69/74.03 67.43/67.75
ReDAD A 88.31/90.15 88.73/89.18 87.49/89.19 77.64/77.07 78.40/76.38 76.04/75.96 77.34/77.21 77.68/76.66 76.45/75.62
D 87.33/88.99 88.74/89.72 87.56/89.63 78.28/76.73 78.27/77.03 76.49/76.02 78.26/76.80 77.89/77.18 76.25/76.04
E 88.03/88.63 88.43/90.48 87.71/87.86 77.83/77.45 78.96/78.02 75.35/73.65 77.75/77.34 78.00/78.10 75.36/73.28
F 87.71/89.54 88.62/89.86 87.63/88.94 77.59/76.50 77.43/77.25 76.34/75.93 77.35/76.33 77.48/77.45 75.47/76.10
G 88.17/90.06 88.72/90.38 87.55/88.85 77.83/77.64 77.76/76.99 76.35/75.71 77.50/77.69 77.33/76.41 76.61/75.73
O 88.27/90.59 88.58/89.91 86.19/87.74 77.08/79.02 76.89/74.85 71.69/71.32 76.78/78.95 76.62/74.93 70.57/70.75
  • 1
    Type of Relation Representation.
  • 2
    Base version.
TABLE VI: LaMEL Results (%) in binary setup (BiomedBERT $\square$: base, $\blacksquare$:large): Each cell (unless cross-disease experiments) shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
@@ IV-D Experiments with Additional Language Models
TABLE VII: LaMEL Results (%) in binary setup utilizing BiomedBERT-base, BioLinkBERT-base, and BioGPT as backbone language model: Each cell shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
Type1 ReDReS ReDAD
BiomedBERT2 BioLinkBERT2 BioGPT BiomedBERT2 BioLinkBERT2 BioGPT
A 90.25/89.43 90.68/89.02 89.43/86.64 86.73/88.75 87.71/89.09 86.14/88.54
D 90.47/88.90 90.35/89.30 88.52/88.17 86.29/88.88 87.35/88.70 85.28/87.85
E 90.61/89.10 90.25/88.08 89.48/87.24 86.03/88.96 87.05/89.36 87.42/87.96
H 89.68/89.24 90.73/89.48 89.21/87.73 87.29/89.91 87.56/88.83 85.55/88.23